Consider the current stock price and two option quotes for the stock of Gone-a-Golf’n, Inc., a firm specializing in high-end golf vacations for the 1%.

Stock            X Call

Price | Exp | Strike | Ask

$10 | Sep |   $11 | $2

$10 | Sep | $8 | $5

Calculate the Intrinsic Value and Time Value of the two options above. Indicate whether each option is in, at, or out of the money.
Answer #1

Intrinsic Value=MAX(Stock Price-Strike Price,0)

1.
=MAX(10-11,0)
=0

2.
=MAX(10-8,0)
=2

Time Value=Premium-Intrinsic Value

1.
=2-0
=2

2.
=5-2
=3

Moneyness:
1.
Out of the money as strike is more than spot price

2.
In the money as strike is less than spot price